OpenStudy (anonymous):

A line passes through (4, 5) and (8, 9). Which equation best represents the line? y = x + 1 y = 1 y = 2x + 1 y = 5x + 8 @AlexandervonHumboldt2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(9-5)/(8-4)=4/4=1 y=x+1

OpenStudy (alexandervonhumboldt2):

A

OpenStudy (anonymous):

thx! :)